# SonarSplat: Novel View Synthesis of Imaging Sonar via Gaussian Splatting

### v2

@@ Abstract
In this paper, we present SonarSplat, a novel Gaussian splatting framework for imaging sonar that demonstrates realistic novel view synthesis and models acoustic streaking phenomena. Our method represents the scene as a set of 3D Gaussians with acoustic reflectance and saturation properties. We develop a novel method to efficiently rasterize Gaussians to produce a range/azimuth image that is faithful to the acoustic image formation model of imaging sonar. In particular, we develop a novel approach to model azimuth streaking in a Gaussian splatting framework. We evaluate SonarSplat using real-world datasets of sonar images collected from an underwater robotic platform in a controlled test tank and in a real-world river environment. Compared to the state-of-the-art, SonarSplat offers improved image synthesis capabilities (+3.2 dB PSNR) and more accurate 3D reconstruction (52% lower Chamfer Distance). We also demonstrate that SonarSplat can be leveraged for azimuth streak removal.
@@ I Introduction
Acoustic sensors, such as imaging sonar, are commonly used for infrastructure inspection, large-area mapping, and target detection in underwater environments [12, 39, 24]. Unlike optical sensors, which are severely range-limited due to water column effects on light propagation, acoustic sensors can capture data at long ranges to provide critical information about subsea environments. Although sonar exhibits many desirable qualities such as longer range, invariance to lighting conditions, and the ability to discern material properties, there exist acoustic phenomena, including elevation ambiguity, azimuth streaking, and multi-path reflections, which make sonar interpretation difficult for operators and computer vision algorithms alike.
@@
Radiance field methods offer a promising approach to addressing these challenges. Neural radiance fields (NeRFs) have demonstrated the potential for high-fidelity data synthesis, denoising, and 3D reconstruction using optical imagery [26, 45, 30, 5]. The benefits of neural rendering have been recognized by the underwater perception community, with prior work exploring neural rendering for 3D object reconstruction using underwater cameras [18, 41] and sonar data [33, 34, 50, 6, 49]. While these innovations are promising, training and processing NeRFs is costly and time-consuming, making deploying them in real-time or on resource-constrained devices difficult. More recently, 3D Gaussian splatting (3DGS) [14] was developed as a faster alternative to NeRF. Gaussian splatting has been leveraged for underwater applications as well [51, 19, 27]. Most relevant to our proposed work, ZSplat proposes a Gaussian splatting framework for RGB-sonar fusion [35]. However, ZSplat leverages the fusion of sonar data to improve the rendering of RGB images and does not focus on enabling high-fidelity data synthesis for sonar imagery or provide evaluation for the quality of rendered sonar images or 3D reconstruction. Thus, there is a clear gap in the literature for a framework capable of efficient and effective sonar image synthesis and 3D reconstruction using Gaussian splatting.
In this paper, we propose SonarSplat (Fig. 1), a novel Gaussian splatting framework for imaging sonar that enables efficient and high quality novel view synthesis, sonar image de-streaking, and 3D reconstruction for underwater robotic applications. To the best of our knowledge, SonarSplat is the first sonar-only method that leverages Gaussian splatting to perform these tasks. Our main contributions are summarized as follows:
  • We develop a novel 3D Gaussian splatting framework for rasterization of range/azimuth sonar images from known and novel viewpoints.
@@
Code and data will be made publicly available on a project website after review.
@@ II Related Work
Figure 2: Overview of SonarSplat. Our method takes as input a sensor pose and an initial set of 3D Gaussians representing the scene. Then, we transform the 3D Gaussians into the sensor’s image space. We splat the reflectivity parameter $\nu_{k}$ to get the unsaturated image $I_{u}$. Additionally, we optimize and splat per-Gaussian azimuth streaking probabilities $p_{k}$. All of the probabilities in a range bin $r_{i}$ are considered in our novel Adaptive Gain Module, which adjusts the receiver gain applied to $I_{u}$. Finally, we produce $\hat{I}$ by multiplying the gain $A$ by $I_{u}$. All parameters are optimized using gradient descent by taking losses with respect to the sonar image pixel values. All sonar images shown are polar (range/azimuth) coordinates.
@@ II-A Radiance Field Methods
NeRFs represent scenes implicitly using neural networks and have been explored for a variety of robotic applications, including novel view synthesis, data generation, Simultaneous Localization and Mapping (SLAM), manipulation, image dehazing, and trajectory optimization [26, 2, 11, 37, 43, 5, 7]. Despite their success as a scene representation, a key challenge for the practical application of NeRFs is the long training time and the slow rendering speeds, primarily due to the computationally expensive volumetric rendering process [26]. A recent development in radiance field methods is Gaussian splatting, which uses more efficient rasterization instead of volumetric rendering to optimize and produce a scene representation, enabling usage in a variety of real-time applications [14, 23, 22, 25].
Despite their impressive performance and rapid adoption, radiance field methods were developed with optical sensors in mind, limiting their direct application to non-visual sensors. Recent developments have integrated the physics-based image formation models of various non-visual sensors, including radar, LIDAR, sonar, thermal cameras, and event cameras into NeRF and Gaussian splatting frameworks for various vision tasks [3, 4, 9, 33, 20]. We are inspired by the faster rendering speeds and high fidelity of Gaussian splatting methods and propose a novel Gaussian splatting framework for imaging sonar.
@@ II-B Perception Challenges for Imaging Sonar
Imaging sonar is a popular underwater sensor for perception tasks like object detection, segmentation, reconstruction, and inspection [33, 8, 12, 44, 28]. However, the sensor is susceptible to multi-path effects, nonlinear noise patterns, and phenomena like azimuth streaking [31]. Azimuth streaking, in particular, can make sonar image understanding difficult by suppressing returns from certain azimuth angles, causing black bars in the sonar image [31]. In addition, the sensor inherently loses information from 3D into 2D when capturing the acoustic returns, causing an effect known as elevation angle ambiguity [46]. This ambiguity has made the successful application of imaging sonar to 3D vision tasks challenging [28, 10, 48].
@@ II-C Novel View Synthesis for Imaging Sonar
Prior work has incorporated the sonar image formation model into signed distance field (SDF) and NeRF formulations [33, 6, 50]. Neusis presents a volumetric rendering framework for dense 3D reconstruction of objects using an imaging sonar [33]. Neusis-NGP is an extension of Neusis that leverages multi-resolution hash encodings to produce a bathymetric heightmap of the environment [50]. Differentiable Space Carving (DSC) proposes a faster method that models echo probabilities to carve out a mesh of the scene [6]. Through our experiments, we show that none of these method are able to perform realistic looking novel view synthesis for imaging sonar, limiting their application to data generation and augmentation. Furthermore, querying these methods is time consuming and limits their application to real-time methods.
To the best of our knowledge, ZSplat is the only Gaussian splatting paper that incorporates imaging sonar data, with demonstrations on underwater data. However, ZSplat oversimplifies the splatting process by not accounting for any sonar-specific phenomena. More importantly, ZSplat only splats the opacities of Gaussians and does not consider material properties like acoustic reflectance. This leads to poor sonar renderings, which limits its application to sonar novel view synthesis. We demonstrate that SonarSplat, which more accurately models sonar image formation, produces more realistic novel view synthesis, restores sonar images to remove azimuth streaking artifacts, and produces 3D reconstructions under a unified framework.
@@ III Technical Approach
Figure 2 shows an overview of SonarSplat, which uses sonar images and their corresponding sensor poses to optimize a scene representation. The scene is represented as a set of 3D Gaussian primitives, each parametrized with a mean, covariance, opacity, acoustic reflectivity and azimuth streaking probability. First, the Gaussians that fall within the view frustum of the sonar are rasterized into a range/azimuth image. Concurrently, we render an azimuth streak probability image using the per-Gaussian azimuth streak probabilities. Based on the relative probabilities in each range bin, we use a novel adaptive gain module to adjust the acoustic returns and capture azimuth streaking effects. Finally, we produce a sonar image with azimuth streaks that can be optimized by taking a loss with respect to the ground truth sonar image. We provide further detail of each of these components in the remainder of this section.
@@ III-A Imaging Sonar Formation Model
To adapt the optical Gaussian splatting framework to accommodate imaging sonar, we must incorporate the acoustic image formation model. Imaging sonar is a time-of-flight sensor that can provide the echo intensities at a given range and azimuth bin. Importantly, imaging sonar is not able to distinguish the elevation angle of a return, leading to an elevation angle ambiguity. This ambiguity makes direct use of imaging sonar difficult for 3D reconstruction and triangulation tasks [47, 10].
A sonar image is defined as $I\in\mathbb{R}^{N_{a}\times N_{r}}$ where $N_{a}$ denotes the number of azimuth bins and $N_{r}$ denotes the number of range bins. In spherical coordinates, the sonar has a vertical field of view ($\phi_{min},\phi_{max}$), a horizontal field of view ($\theta_{min},\theta_{max}$), and a maximum range, $R_{max}$. Finally, taking $\frac{R_{max}}{N_{r}}$ gives us the range resolution, $\epsilon_{r}$, and taking $\frac{\theta_{max}-\theta{min}}{N_{a}}$ gives us the azimuthal resolution, $\epsilon_{a}$.
@@
where $E_{e}$ is the intensity of emitted sound, $r$ is the range of the point, $T(r,\theta,\phi)$ is the transmittance term, and $\sigma(r,\theta,\phi)$ is the density of the point. Note that this transmittance term is equivalent to the term used in volumetric rendering, serving as the probability that a ray travels to point $(r,\theta,\phi)$ without hitting another particle.
III-B Range/Azimuth Splatting Model
To evaluate this integral, prior methods use sampling and quadrature [33, 34, 6]. However, motivated by the increased efficiency of splatting techniques, we explore a point-based approach adapted to the sonar image formation model. To this end, we represent the scene with a set of 3D Gaussian primitives $\mathcal{G}=\{G_{k}\}_{k=1}^{N}$. We augment this parametrization with acoustic-specific variables: acoustic reflectivity $\nu_{k}$ and azimuth streak probability $p_{k}$. Formally, these Gaussians are parametrized with a mean $\mu_{k}\in\mathbb{R}^{3}$, scale $s_{k}\in\mathbb{R}^{3}$, orientation $q_{k}\in\mathbb{H}$, opacity $o_{k}\in\mathbb{R}^{+}$, acoustic reflectivity $\nu_{k}\in\mathbb{R}^{+}$, and azimuth streak probability $p_{k}\in\mathbb{R}^{+}$:
$G_{k}=\{\mu_{k},s_{k},q_{k},o_{k},\nu_{k},p_{k}\}$ (2)
where the covariance $\Sigma_{k}$ is computed as in [14].
To render sonar images, we first transform the means and covariances of the Gaussians into spherical coordinates. Let $\mu_{k}=\langle x_{k},y_{k},z_{k}\rangle~{}\text{and}~{}\Sigma_{k}$ be the mean and covariance in the sensor’s frame, respectively. The transformation to spherical coordinates is as follows [15]:
$\mu_{s}=\begin{bmatrix}r_{k}\\ \theta_{k}\\ \phi_{k}\end{bmatrix}=\begin{bmatrix}||\mu_{k}||\\ \arctan(y_{k},x_{k})\\ \arctan(z_{k},\sqrt{x_{k}^{2}+y_{k}^{2}})\\ \end{bmatrix}$ (3)
@@ III-B Range/Azimuth Splatting Model
$J_{s}=\begin{bmatrix}\frac{x_{k}}{\lVert\mu_{k}\rVert}&\frac{y_{k}}{\lVert\mu_{k}\rVert}&\frac{z_{k}}{\lVert\mu_{k}\rVert}\\[4.30554pt] \frac{-y}{x^{2}+y^{2}}&\frac{x}{x^{2}+y^{2}}&0\\[4.30554pt] \frac{-x_{k}\cdot z_{k}}{||\mu_{k}||^{2}\sqrt{x^{2}+y^{2}}}&\frac{-y_{k}\cdot z_{k}}{||\mu_{k}||^{2}\sqrt{x^{2}+y^{2}}}&\frac{\sqrt{x^{2}+y^{2}}}{\lVert\mu_{k}\rVert^{2}}\end{bmatrix}$ (4)
@@
and then we perform the transformation
@@
To find the intensity of acoustic returns at a given range/azimuth bin ($r_{i},\theta_{j}$), we consider $\mathcal{G}_{i,j}\subset\mathcal{G}$, the set of Gaussians transformed to range/azimuth image space that overlap with the range/azimuth bin ($r_{i},\theta_{j}$). The rasterization equation to obtain the sonar image for SonarSplat is similar to that of [14], with a few key differences. SonarSplat’s rasterization equation is given by:
$I_{u}(r_{i},\theta_{j})=\frac{1}{R_{k}}\sum_{G_{k}^{\prime}\in\mathcal{G}_{i,j}}\nu_{k}~{}o_{k}~{}T_{k}~{}G_{k}^{\prime}(r_{i},\theta_{j})$ (8)
where $R_{k}$ is the range from the sensor, $T_{k}$ is the transmittance from the sensor origin to the Gaussian along the range, and $G^{\prime}_{k}$ is the Gaussian evaluated in range/azimuth space. The subscript indicates that $I_{u}$ is the unsaturated image, meaning that it does not account for the azimuth streaking that causes saturation across the image. We observe that the acoustic reflectance of a set of points depends on material properties and viewing angle, so $\nu_{k}$ uses spherical harmonics to encode view-dependent effects. In practice, our sonar images are gain adjusted, so $\frac{1}{R_{k}}$ is dropped.

### v3

@@ Abstract
In this paper, we present SonarSplat, a novel Gaussian splatting framework for imaging sonar that demonstrates realistic novel view synthesis and models acoustic streaking phenomena. Our method represents the scene as a set of 3D Gaussians with acoustic reflectance and saturation properties. We develop a novel method to efficiently rasterize Gaussians to produce a range/azimuth image that is faithful to the acoustic image formation model of imaging sonar. In particular, we develop a novel approach to model azimuth streaking in a Gaussian splatting framework. We evaluate SonarSplat using real-world datasets of sonar images collected from an underwater robotic platform in a controlled test tank and in a real-world river environment. Compared to the state-of-the-art, SonarSplat offers improved image synthesis capabilities (+3.2 dB PSNR) and more accurate 3D reconstruction (77% lower Chamfer Distance). We also demonstrate that SonarSplat can be leveraged for azimuth streak removal.
@@ I Introduction
Acoustic sensors, such as imaging sonar, are commonly used for infrastructure inspection, large-area mapping, and target detection in underwater environments [1, 2, 3]. While optical sensors are severely range-limited due to water column effects on light propagation, acoustic sensors can capture data at long ranges to provide critical information about subsea environments. However, acoustic phenomena like elevation ambiguity, azimuth streaking, and multi-path reflections make sonar interpretation difficult for both operators and computer vision algorithms.
@@
Radiance field methods offer a promising approach to addressing these challenges. Neural radiance fields (NeRFs) have demonstrated the potential for high-fidelity data synthesis, denoising, and 3D reconstruction using optical imagery [4, 5, 6, 7]. The benefits of neural rendering have been recognized by the underwater perception community, with prior work exploring neural rendering for 3D object reconstruction using underwater cameras [8, 9] and sonar data [10, 11, 12, 13, 14]. While these innovations are promising, training and processing NeRFs is costly and time-consuming, making deploying them in real-time or on resource-constrained devices difficult. More recently, 3D Gaussian splatting (3DGS) [15] was developed as a faster alternative to NeRF. Gaussian splatting has been leveraged for underwater applications as well [16, 17, 18]. Most relevant to our proposed work, ZSplat proposes a Gaussian splatting framework for RGB-sonar fusion [19]. However, ZSplat leverages the fusion of sonar data to improve the rendering of RGB images and does not focus on enabling high-fidelity data synthesis for sonar imagery or providing evaluation for the quality of rendered sonar images or 3D reconstruction. Thus, there is a clear gap in the literature for a framework capable of efficient and effective sonar image synthesis and 3D reconstruction using Gaussian splatting.
In this paper, we propose SonarSplat (Fig. 1), a novel Gaussian splatting framework for imaging sonar that enables efficient and high-quality novel view synthesis, sonar image de-streaking, and 3D reconstruction for underwater robotic applications. To the best of our knowledge, SonarSplat is the first sonar-only method that leverages Gaussian splatting to perform these tasks. Our main contributions are as follows:
  • We develop a novel 3D Gaussian splatting framework for rasterization of range/azimuth sonar images from known and arbitrary viewpoints.
@@
For further information, please visit our project website at https://umfieldrobotics.github.io/sonarsplat3D/.
@@ II Related Work
Figure 2: Overview of SonarSplat. Our method takes as input a sensor pose and an initial set of 3D Gaussians representing the scene. Then, we transform the 3D Gaussians into the sensor’s image space. We splat the reflectivity parameter $\nu_{k}$ to get the unsaturated image $I_{u}$. Additionally, we optimize and splat per-Gaussian azimuth streaking probabilities $p_{k}$. All of the probabilities in a range bin $r_{i}$ are considered in our novel Adaptive Gain Module, which adjusts the gain AA applied to $I_{u}$. Finally, we produce $\hat{I}$ by multiplying the gain AA by $I_{u}$. All parameters are optimized using gradient descent by taking losses with respect to the sonar image pixel values. All sonar images shown are polar (range/azimuth) coordinates.
@@ II-A Radiance Field Methods
NeRFs represent scenes implicitly using neural networks and have been explored for a variety of robotic applications, including novel view synthesis [4], data generation [20], Simultaneous Localization and Mapping (SLAM) [21, 22], manipulation [23], and image dehazing [7]. Despite their success as a scene representation, a key challenge for the practical application of NeRFs is the long training time and the slow rendering speeds, primarily due to the computationally expensive volumetric rendering process [4]. A recent development in radiance field methods is Gaussian splatting, which uses more efficient rasterization instead of volumetric rendering to produce a scene representation. This has led to the use of radiance field methods in a variety of real-time applications [15, 24, 25].
Despite their impressive performance and rapid adoption, radiance field methods were developed with optical sensors in mind, limiting their direct application to non-visual sensors. Recent developments have integrated non-visual sensors, including radar, LIDAR, sonar, thermal cameras, and event cameras into NeRF and Gaussian splatting frameworks for various vision tasks [26, 27, 21, 28, 10, 29]. We are inspired by the faster rendering speeds and high fidelity of Gaussian splatting methods and propose a novel Gaussian splatting framework for imaging sonar.
@@ II-B Perception Challenges for Imaging Sonar
Imaging sonar is a popular underwater sensor for perception tasks like object detection, segmentation, reconstruction, and inspection [10, 30, 1, 31, 32]. However, the sensor is susceptible to multi-path effects, nonlinear noise patterns, and phenomena like azimuth streaking [33]. Azimuth streaking, in particular, can make sonar image understanding difficult by suppressing returns from specific ranges, causing black bars in the sonar image [33]. In addition, the sensor inherently loses information from 3D into 2D when capturing the acoustic returns, causing an effect known as elevation angle ambiguity [34]. This ambiguity has made the successful application of imaging sonar to 3D vision tasks challenging [32, 35, 36].
@@ II-C Novel View Synthesis for Imaging Sonar
Prior work has incorporated the sonar image formation model into signed distance field (SDF) and NeRF formulations [10, 13, 12]. Neusis presents a volumetric rendering framework for dense 3D reconstruction of objects using an imaging sonar [10]. Neusis-NGP is an extension of Neusis that leverages multi-resolution hash encodings to produce a bathymetric heightmap of the environment [12]. Differentiable Space Carving (DSC) proposes a faster method that models echo probabilities to carve out a mesh of the scene [13]. Through our experiments, we demonstrate that these methods render less realistic sonar imagery compared to SonarSplat according to quantitative metrics including PSNR, SSIM, and LPIPS. This limits their use in data generation and augmentation. Additionally, our experiments indicate that querying these methods is time consuming, which limits their use in real-time applications.
To the best of our knowledge, ZSplat is the only Gaussian splatting paper that incorporates imaging sonar data, with demonstrations on underwater data. However, ZSplat operates with a simplified sonar model that does not model several important sonar-specific phenomena. More importantly, ZSplat only splats the opacities of Gaussians and does not consider material properties like acoustic reflectance. This leads to poor sonar renderings, which limits its application in sonar novel view synthesis. We demonstrate that SonarSplat, which more accurately models sonar image formation, produces more realistic novel view synthesis, restores sonar images to remove azimuth streaking artifacts, and produces 3D reconstructions under a unified framework.
@@ III Technical Approach
Figure 2 shows an overview of SonarSplat, which uses sonar images and their corresponding sensor poses to optimize a scene representation. The scene is represented as a set of 3D Gaussian primitives, each parametrized with a mean, covariance, opacity, acoustic reflectivity, and azimuth streaking probability. First, the Gaussians that fall within the view frustum of the sonar are rasterized into a range/azimuth image. Concurrently, we render an azimuth streak probability image using the per-Gaussian azimuth streak probabilities. Based on the relative probabilities in each range bin, we use a novel adaptive gain module to adjust the acoustic returns and capture azimuth streaking effects. Finally, we produce a sonar image with azimuth streaks that can be optimized by taking a loss with respect to the ground truth sonar image. We provide further details of each of these components in the remainder of this section.
@@ III-A Imaging Sonar Formation Model
To adapt the optical Gaussian splatting framework for imaging sonar, we must incorporate the acoustic image formation model into the rasterization process. Imaging sonar is a time-of-flight sensor that provides the echo intensities at a given range and azimuth bin.
A sonar image is defined as $I\in\mathbb{R}^{N_{a}\times N_{r}}$ where $N_{a}$ denotes the number of azimuth bins and $N_{r}$ denotes the number of range bins. In spherical coordinates, the sonar has a vertical field of view ($\phi_{min},\phi_{max}$), a horizontal field of view ($\theta_{min},\theta_{max}$), and a maximum range, $R_{max}$. Finally, $\frac{R_{max}}{N_{r}}$ is the range resolution, $\epsilon_{r}$, and $\frac{\theta_{max}-\theta_{min}}{N_{a}}$ is the azimuthal resolution, $\epsilon_{a}$.
@@
$I(r_{i},\theta_{j})=\int_{\phi_{min}}^{\phi_{max}}\int_{\theta_{j}-\epsilon_{a}}^{\theta_{j}+\epsilon_{a}}\int_{r_{i}-\epsilon_{r}}^{r_{i}+\epsilon_{r}}\frac{E_{e}}{r}T(r,\theta,\phi)\sigma(r,\theta,\phi)drd\theta d\phi$
where $E_{e}$ is the intensity of emitted sound, $r$ is the range of the point, $T(r,\theta,\phi)$ is the transmittance term, and $\sigma(r,\theta,\phi)$ is the density of the point. $T(r,\theta,\phi)$ is the probability that a ray travels to point $(r,\theta,\phi)$ without hitting another particle.
III-B Range/Azimuth Splatting Model
To evaluate this integral, prior methods use sampling and quadrature [10, 11, 13]. However, motivated by the increased efficiency of splatting techniques, we explore a point-based approach adapted to the sonar image formation model. To this end, we represent the scene with a set of 3D Gaussian primitives $\mathcal{G}=\{G_{k}\}_{k=1}^{N}$. Formally, these Gaussians are parametrized with a mean $\mu_{k}\in\mathbb{R}^{3}$, scale $s_{k}\in\mathbb{R}^{3}$, orientation $q_{k}\in\mathbb{H}$, opacity $o_{k}\in\mathbb{R}^{+}$, and azimuth streak probability $p_{k}\in\mathbb{R}^{+}$. We take inspiration from other Time-of-Flight novel view synthesis work and use spherical harmonics coefficients $sh_{k}$ to model view-dependent acoustic reflectivity $\nu_{k}$ [28]. Specifically, $\nu_{k}=SH(sh_{k},\vec{v})$, where $\vec{v}$ is the viewing direction.
$G_{k}=(\mu_{k},s_{k},q_{k},o_{k},sh_{k},p_{k})$ (2)
where the covariance $\Sigma_{k}$ $R_{k}S_{k}S_{k}^{T}R_{k}^{T}$ for rotation matrix $R_{k}$ given by quaternion $q_{k}$ and scaling matrix $S_{k}$ given by $s_{k}$.
To render sonar images, we first transform the means and covariances of the Gaussians into spherical coordinates. Let $\mu_{k}=[x_{k},y_{k},z_{k}]^{T}~\text{and}~\Sigma_{k}$ be the mean and covariance in the sensor’s frame, respectively. The transformation to spherical coordinates is as follows [37]:
@@ III-B Range/Azimuth Splatting Model
[…]
@@
and then we perform the transformation:
@@
To find the intensity of acoustic returns at a given range/azimuth bin $r_{i},\theta_{j}]$, we consider $\mathcal{G}_{i,j}\subset\mathcal{G}$, the set of Gaussians transformed to range/azimuth image space that overlap with the range/azimuth bin $r_{i},\theta_{j}]$. The rasterization equation to obtain the sonar image for SonarSplat is similar to that of [15], with a few key differences. Specifically, we use single-channel acoustic reflectance $\nu$ instead of per-channel color $c$ and we form the image by splatting in the range/azimuth plane using $r,\theta]$ coordinates instead of traditional camera coordinates. SonarSplat’s rasterization equation is given by:
$I_{u}(r_{i},\theta_{j})=\sum_{G_{k}\in\mathcal{G}_{i,j}}\frac{1}{r_{k}}\nu_{k}~o_{k}~\overline{T}_{k}~f_{k}([r_{i},\theta_{j}])$ (8)
$f_{k}(x)=\text{exp}(-\frac{1}{2}\lVert x-\mu^{\prime}_{k}\rVert^{2}_{\Sigma^{\prime}_{k}})$ (9)
$\overline{T}_{k}=\frac{1}{|\mathcal{P}_{k}|}\sum_{p\in\mathcal{P}_{k}}T^{p}_{k}$ (10)
where $r_{k}$ is the range of $G_{k}$ from the sensor, $f_{k}(x)$ is the evaluation of Gaussian $G_{k}$ at range/azimuth point $r_{i},\theta_{j}]$ with the norm as the Mahalanobis distance, $\overline{T}_{k}$ is the average transmittance of the Gaussian from the sensor origin to the Gaussian along the range. To compute $\overline{T}_{k}$, 3D Gaussians are splatted onto the azimuth/elevation plane to model occlusions in the range dimension. We define $\mathcal{P}_{k}$ as the set of azimuth/elevation pixels that the Gaussian $G_{k}$ overlaps and we calculate $T^{p}_{k}$ as the transmittance of $G_{k}$ at pixel $p$ according to [15]. The subscript indicates that $I_{u}$ is the unsaturated image, meaning that it does not account for the azimuth streaking that causes saturation across the image. Our sonar images are gain adjusted, so $\frac{1}{r_{k}}$ is dropped.